# Stability phenomena in Deligne-Mumford compactifications via Morse theory

### v1

@@ Abstract.
We study the rational homology of the Deligne-Mumford compactification $\overline{\mathcal{M}}_{g,n}$ of the moduli space of stable curves via a family of Morse functions, the $\operatorname{sys_{T}}$ functions, which encode geometric information about short geodesics on hyperbolic surfaces. Exploiting the Morse-theoretic properties of $\operatorname{sys_{T}}$, including the existence of an index gap and the behavior of critical points near boundary strata, we prove that in low degrees the homology of $\overline{\mathcal{M}}_{g,n}$ is supported entirely on the boundary $\partial\mathcal{M}_{g,n}$.
Furthermore, we establish finite generation and stability phenomena for the rational homology across all genera and numbers of marked points. Using stable graphs and explicit attaching maps, we show that for each fixed degree $k$, a finite set of critical points generates all $k$-th homology classes via attaching stable graphs of stratum dimension 0. This result recovers previously known stability in the genus direction, such as Tosteson’s theorem, and provides a concrete, geometric construction of homology generators. Our approach unifies combinatorial, geometric, and Morse-theoretic techniques to give a comprehensive picture of the low-degree and stable homology of $\overline{\mathcal{M}}_{g,n}$.
@@ 1. Introduction
The moduli space of stable curves $\overline{\mathcal{M}}_{g,n}$ is a central object in algebraic geometry and low-dimensional topology. Its rich structure combines complex analytic, hyperbolic, and combinatorial perspectives, making it an ideal testing ground for techniques from Morse theory, Teichmüller theory, and homological stability. Understanding the topology of $\overline{\mathcal{M}}_{g,n}$, particularly its rational homology in various degrees, remains a fundamental question with deep connections to mapping class groups and representation stability.
In this paper, we study the low-degree homology and stability phenomena of $\overline{\mathcal{M}}_{g,n}$ via the $\operatorname{sys_{T}}$ functions, a family of $C^{2}$-Morse functions defined on the moduli space. These functions, inspired by systolic geometry, encode geometric information about short geodesics on hyperbolic surfaces and extend $C^{2}$-continuously to the Deligne-Mumford compactification. Their Morse-theoretic properties, including the existence of an index gap and the behavior of critical points near boundary strata, provide a powerful tool for analyzing homology in a combinatorial and geometric fashion.
Our first main result concerns the low-degree homology of $\overline{\mathcal{M}}_{g,n}$. By exploiting the $\operatorname{sys_{T}}$-Morse theory and the structure of the Deligne-Mumford boundary, we prove:
@@
$i\colon\partial\mathcal{M}_{g,n}\to\overline{\mathcal{M}}_{g,n}$
@@
$i_{*}\colon H_{k}(\partial\mathcal{M}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
This theorem implies that, in the stable range, low-degree homology classes are determined by boundary strata, a phenomenon revealed by the classical homological stability results of Harer [harer1985stability, harer1986virtual].
Our second main contribution addresses stability phenomena across the moduli spaces $\overline{\mathcal{M}}_{g,n}$ as $(g,n)$ varies. Using the combinatorial framework of stable graphs and the $\operatorname{sys_{T}}$-Morse theory, we establish that rational homology in each fixed degree is finitely generated and can be constructed via explicit attaching maps of low-dimensional stable graphs:
@@
via attaching a stable graph of stratum dimension 0.
This result not only provides a concrete description of generators for homology across all $(g,n)$ but also recovers previously known stability results in the genus direction, such as Tosteson’s theorem [tosteson2021stability]. In particular, our approach gives a geometric and Morse-theoretic perspective on the phenomena of homological stabilization and the generation of homology by boundary strata.
The structure of the paper is as follows. In Section 2, we review stable graphs, strata of $\overline{\mathcal{M}}_{g,n}$, and the definition of the $\operatorname{sys_{T}}$ functions. Section 3 discusses the Weil–Petersson metric and the construction of gradient systems compatible with the $\operatorname{sys_{T}}$ functions. Section 4 proves the main low-degree homology isomorphism, and Section 5 establishes the finite generation and stability phenomena for the rational homology of $\overline{\mathcal{M}}_{g,n}$.
Throughout, we work over the rational coefficients unless otherwise specified. The techniques developed here combine geometric, combinatorial, and Morse-theoretic ideas, offering a unified approach to understanding the topology of moduli spaces of curves.
@@ 1.1. Acknowledgements:
The author is grateful to Louis Hainaut, Siddarth Kannan, Jeremy Kahn, and Philip Tosteson for valuable conversations and insights that contributed to the development of this work.
@@ 2. Stable graphs and the $\operatorname{sys_{T}}$ functions
We review standard definitions used to study Deligne–Mumford compactifications and discuss the construction and properties of the $\operatorname{sys_{T}}$ functions.
2.1. Stable graph and Deligne-Mumford compactification
@@ 2.2. The Morse functions
There is a canonical correspondence between algebraic structures and hyperbolic structures on a topological surface. Thus, the moduli space can be defined as the quotient of the space of hyperbolic metrics by the action of orientation-preserving diffeomorphisms.
@@
where s.c.g. stands for simple closed geodesic. The definition of $\operatorname{sys_{T}}$ on the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$ is by continuous extension.
@@
A pointed subspace of $\overline{\mathcal{M}}_{g,n}$ is called stratifically stable if the subspace is contained in the relative Deligne-Mumford compactification of the stratum the point lies in. The following theorem establishes the Morse property of $\operatorname{sys_{T}}$, which is foundational to subsequent Morse homology results.
@@
  • The $\operatorname{sys_{T}}$ function is $C^{2}$-Morse on the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$.
@@
3. Subtleties in using the Weil–Petersson metric for Morse theory
@@ 3.1. Weil–Petersson metric near the boundary
The Weil–Petersson metric, induced by the natural $L^{2}$-pairing on the space of (0,2)- or (-1,1)-tensors, is known to be incomplete on Teichmüller space. Consequently, when extended to the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$, it defines an orbifold Riemannian metric that is singular on the boundary $\partial\mathcal{M}_{g,n}$. Wolpert [wolpert2009extension] computed the expansion of the Weil–Petersson connection near the boundary of $\overline{\mathcal{M}}_{g,n}$.
@@ 4. Low degree homology
$i\colon\partial\mathcal{M}_{g,n}\to\overline{\mathcal{M}}_{g,n}$
@@
$i_{*}\colon H_{k}(\partial\mathcal{M}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
@@
  • Near a boundary stratum $\mathcal{S}_{S}$, where $S=\{\gamma_{i}\}$ is a set of mutually disjoint simple closed geodesics, we have
@@
Let $M$ be a finite manifold cover of $\overline{\mathcal{M}}_{g,n}$, and let $B\subset M$ denote the preimage of $\partial\mathcal{M}_{g,n}$. By abuse of notation, we denote the lift of the $\operatorname{sys_{T}}$ function to $M$ also by $\operatorname{sys_{T}}$.
@@
$B=\operatorname{sys}^{-1}(0)\cong M^{\leq\epsilon}\cong\operatorname{sys_{T}}^{-1}([0,\delta]).$
@@
$i_{*}\colon H_{k}(\partial\mathcal{M}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
Taking the long exact sequence of the pair $(\overline{\mathcal{M}}_{g,n},\partial\mathcal{M}_{g,n})$, we deduce:
@@
$H^{6g-6+2n-k}(\mathcal{M}_{g,n};\mathbb{Q})\cong H^{c}_{k}(\mathcal{M}_{g,n};\mathbb{Q})\cong H_{k}(\overline{\mathcal{M}}_{g,n},\partial\mathcal{M}_{g,n};\mathbb{Q})=0.$
@@
This result is equivalent to Harer’s homological stability [harer1985stability, harer1986virtual] in the stable range, although here we only consider a weaker statement for sufficiently large $g+n$.
@@ 5. Stability Phenomena
$\mathcal{S}_{g,n}([k])\subset\partial\mathcal{M}_{g,n}.$
@@
In other words, when $\operatorname{sys_{T}}$-Morse theory applies, the “complexity” of $k$-th rational homology elements stabilizes as either $g$ or $n$ increases.
@@
By composing attaching maps, one can attach a stable graph of stratum dimension 0 (dual to a maximally noded curve) to stable graphs with marked points. Morse-theoretically, attaching copies of thrice-marked $\mathbb{P}^{1}$ preserves criticality and Morse index, inducing equivariant homomorphisms on rational homology and Morse homology.
@@
via attaching a stable graph of stratum dimension 0.
@@
Theorem 5.6 recovers the stability in $g$ previously established by Tosteson [tosteson2021stability], noting that a Whitehead move on a stable graph preserves homology up to a nonzero factor.
Theorem 5.8 ([tosteson2021stability]).
Let $i,g\in\mathbb{N}$. There exists $N\in\mathbb{N}$ such that the natural map of $\text{FS}^{\text{op}}$-modules
$\text{Ind}_{N}\text{Res}_{N}H_{i}(\mathcal{M}_{g,-},\mathbb{Q})\to H_{i}(\mathcal{M}_{g,-},\mathbb{Q})$
is an isomorphism. In particular, any presentation of the $\text{FS}_{N}^{\text{op}}$-module $\text{Res}_{N}H_{i}(\mathcal{M}_{g,-},\mathbb{Q})$ gives a presentation of the $\text{FS}^{\text{op}}$-module $H_{i}(\mathcal{M}_{g,-},\mathbb{Q})$.

### v2

# Stability phenomena in Deligne–Mumford compactifications via Morse theory

@@ Abstract.
We study the rational homology of the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$ of the moduli space of stable curves via a family of Morse functions, namely the $\text{sys}_{T}$ functions. Exploiting the geometric and Morse properties of $\text{sys}_{T}$, including the existence of an index gap and additivity of the Morse index upon gluing maps, we reprove that in low degrees the homology of $\overline{\mathcal{M}}_{g,n}$ is supported entirely on the boundary $\partial\overline{\mathcal{M}}_{g,n}$, providing a geometric perspective complementary to Harer’s classical result on the virtual cohomological dimension.
Furthermore, we establish finite generation and stability phenomena for the rational homology across all genera and numbers of marked points. We show that for each degree $k$, a finite set of homology elements generates all $k$-th homology classes via attaching copies of thrice-marked $\mathbb{P}^{1}$. This result also recovers previously known stability in the number of marked points, such as Tosteson’s theorem.
@@ 1. Introduction
The moduli space of stable curves $\overline{\mathcal{M}}_{g,n}$ is a central object in algebraic geometry and geometric topology. Its rich structure combines complex analytic, hyperbolic, and combinatorial perspectives. Understanding the topology of $\overline{\mathcal{M}}_{g,n}$, particularly its rational homology in various degrees, remains a fundamental question.
In this paper, we study the low-degree homology and stability phenomena of $\overline{\mathcal{M}}_{g,n}$ via the $\operatorname{sys_{T}}$ functions, a family of $C^{2}$-Morse functions. These functions, inspired by systolic geometry, encode geometric information about closed geodesics on hyperbolic surfaces and continuously extend to the Deligne–Mumford compactification. The $\operatorname{sys_{T}}$-Morse theoretic techniques combine geometric, combinatorial, and Morse-theoretic ideas, offering a novel approach to understanding the topology of moduli spaces of curves.
Our first result concerns the low-degree homology of $\overline{\mathcal{M}}_{g,n}$. By exploiting the $\operatorname{sys_{T}}$-Morse theory and the structure of the Deligne–Mumford boundary, we reprove:
@@
$i\colon\partial\overline{\mathcal{M}}_{g,n}\to\overline{\mathcal{M}}_{g,n}$
@@
$i_{*}\colon H_{k}(\partial\overline{\mathcal{M}}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
This theorem implies that, in the stable range, low-degree homology classes are determined by boundary strata, a phenomenon revealed by Harer’s theorem on the virtual cohomological dimension of $\mathcal{M}_{g,n}$ in [harer1986virtual].
Our second main result addresses stability phenomena across the moduli spaces $\overline{\mathcal{M}}_{g,n}$ as $(g,n)$ varies. Let $\operatorname{ind}(g,n)$ be the lowest Morse index of $\operatorname{sys_{T}}$-critical points in $\mathcal{M}_{g,n}$. Using the combinatorial framework of stable graphs and the $\operatorname{sys_{T}}$-Morse theory, we show:
@@
via attaching thrice-marked $\mathbb{P}^{1}$’s. Here, the stable range is given by $\operatorname{ind}(g,n)>k$.
This result provides a concrete description of the generation of homology by boundary strata. It recovers previously known stability results in $n$, the number of marked points, such as Tosteson’s theorem [tosteson2021stability].
The structure of the paper is as follows. In Section 2, we review stable graphs, strata of $\overline{\mathcal{M}}_{g,n}$, and the definition of the $\operatorname{sys_{T}}$ functions. Section 3 discusses the Weil–Petersson metric and potential substitutes for the $\operatorname{sys_{T}}$ functions. Section 4 proves the low-degree homology isomorphism, and Section 5 establishes the finite generation and stability phenomena for the rational homology of $\overline{\mathcal{M}}_{g,n}$. Finally, in Section 6, we relate our results to previous work.
@@ 1.1. Acknowledgements:
The author is grateful to Louis Hainaut, Siddarth Kannan, Jeremy Kahn, and Philip Tosteson for valuable conversations and insights that contributed to the development of this work. The author thanks Oscar Randal-Williams for thoughtful comments and suggestions.
@@ 2. Stable graphs and the $\operatorname{sys_{T}}$ functions
We review standard definitions used to study Deligne–Mumford compactifications and discuss the construction and geometric and Morse properties of the $\operatorname{sys_{T}}$ functions.
2.1. Stable graph and Deligne–Mumford compactification
@@ 2.2. The Morse functions
There is a canonical correspondence between algebraic structures and hyperbolic structures on a topological real surface. Thus, the moduli space $\mathcal{M}_{g,n}$ can be defined as the quotient of the space of hyperbolic metrics by the action of orientation-preserving diffeomorphisms.
@@
where s.c.g. stands for simple closed geodesic. The definition of it on the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$ is by continuous extension.
@@
Definition 2.4.
For a set $B=\{\beta_{1},\dots,\beta_{k}\}$ of pairwise disjoint essential closed geodesics on a given hyperbolic surface, let $D_{B}\subset\overline{\mathcal{M}}_{g,n}$ be the stratum where the curves in $B$ are pinched. Extend $B$ to a pants decomposition and consider the standard Fenchel–Nielsen coordinates $(l_{\gamma},\tau_{\gamma})_{\gamma}$. For $X\in D_{B}$, define the local chart
$(l_{\gamma}^{\chi(\gamma)},\tau_{\gamma})_{\gamma}/\text{Aut}(X),$
where $\chi(\gamma)=\frac{1}{2}$ if $\gamma\in B$ and 1 otherwise. This endows $\overline{\mathcal{M}}_{g,n}$ with the the root-geodesic-length smooth orbifold structure on $\overline{\mathcal{M}}_{g,n}$.
A pointed subspace of $\overline{\mathcal{M}}_{g,n}$ is called stratifically stable if the subspace is contained in the relative Deligne–Mumford compactification of the stratum the point lies in. The following theorem establishes the Morse property of $\operatorname{sys_{T}}$, which is foundational to subsequent Morse homology results.
@@
  • The $\operatorname{sys_{T}}$ function is $C^{2}$-Morse on the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$, equipped with the root-geodesic-length smooth orbifold structure.
@@
Remark 2.8.
At present, this Morse-theoretic approach does not yield optimal numerical results for the (co)homology, primarily because relatively little is known about the following aspects:
(1) The lower bound on $\operatorname{ind}(g,n)$ given above is not sharp. As shown in [bourque2020hyperbolic], a linear bound cannot be achieved, in contrast with many results such as the stable range for the homology of $\mathcal{M}_{g,n}$. Nevertheless, one might still expect a logarithmic bound.
(2) In general, the number of critical points exceeds the dimension of the homology in the corresponding degree. When a Riemannian metric is chosen, the critical points interact with each other via gradient flow lines, and the correspondence between critical points and homology generators remains largely unexplored.
3. Subtleties in using the Weil–Petersson metric
@@ 3.1. Weil–Petersson metric near the boundary
The Weil–Petersson metric, induced by the natural $L^{2}$-pairing on the space of (0,2)- or (-1,1)-tensors, is known to be incomplete on Teichmüller space. Consequently, when extended to the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,n}$, it defines an orbifold Riemannian metric that is singular on the boundary $\partial\overline{\mathcal{M}}_{g,n}$. Wolpert [wolpert2009extension] computed the expansion of the Weil–Petersson connection near the boundary of $\overline{\mathcal{M}}_{g,n}$.
@@ 4. Low degree homology
$i\colon\partial\overline{\mathcal{M}}_{g,n}\to\overline{\mathcal{M}}_{g,n}$
@@
$i_{*}\colon H_{k}(\partial\overline{\mathcal{M}}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
@@
  • Near a boundary stratum $D_{S}$, where $S=\{\gamma_{i}\}$ is a set of mutually disjoint simple closed geodesics, we have
@@
Let $M$ be a finite manifold cover of $\overline{\mathcal{M}}_{g,n}$, and let $B\subset M$ denote the preimage of $\partial\overline{\mathcal{M}}_{g,n}$. By abuse of notation, we denote the lift of the $\operatorname{sys_{T}}$ function to $M$ also by $\operatorname{sys_{T}}$.
@@
$B\times[0,\delta]=\operatorname{sys}^{-1}(0)\times[0,\delta]\cong M^{\leq\epsilon}\cong\operatorname{sys_{T}}^{-1}([0,\delta]).$
@@
$i_{*}\colon H_{k}(\partial\overline{\mathcal{M}}_{g,n};\mathbb{Q})\xrightarrow{\cong}H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q}).$
Let $H^{\text{BM}}_{*}$ denote the Borel–Moore homology. Taking the long exact sequence of the pair $(\overline{\mathcal{M}}_{g,n},\partial\overline{\mathcal{M}}_{g,n})$, we deduce:
@@
$H^{6g-6+2n-k}(\mathcal{M}_{g,n};\mathbb{Q})\cong H^{\text{BM}}_{k}(\mathcal{M}_{g,n};\mathbb{Q})\cong H_{k}(\overline{\mathcal{M}}_{g,n},\partial\overline{\mathcal{M}}_{g,n};\mathbb{Q})=0.$
@@
This result is equivalent to Harer’s theorem on the virtual cohomological dimension [harer1986virtual] in the stable range, although here we only consider a weaker statement for sufficiently large $g+n$.
@@ 5. Stability Phenomena
$\mathcal{S}_{g,n}([k])\subset\partial\overline{\mathcal{M}}_{g,n}.$
@@
In other words, when $\operatorname{sys_{T}}$-Morse theory applies, the ‘complexity’ of $k$-th rational homology elements stabilizes as either $g$ or $n$ increases.
@@
By composing attaching maps, one can attach a stable graph of stratum dimension 0 (dual to a maximally noded curve) to stable graphs with marked points. Morse-theoretically, attaching copies of thrice-marked $\mathbb{P}^{1}$ preserves criticality and the Morse index, inducing equivariant homomorphisms on rational homology and Morse homology.
@@
via attaching thrice-marked $\mathbb{P}^{1}$’s. Here, the stable range is given by $\operatorname{ind}(g,n)>k$.
6. Relation to previous work
6.1. Tosteson’s stability theorem in $n$
Tosteson [tosteson2021stability] studies the asymptotic behavior of the homology of $\overline{\mathcal{M}}_{g,n}$ using $\mathbf{FS}^{\text{op}}$-modules.
Let $\mathbf{FS}$ be the category of finite sets and surjections. An $\mathbf{FS}^{\text{op}}$-module is a functor from $\mathbf{FS}^{\text{op}}$ to the category of vector spaces, denoted $n\mapsto V_{n}$, and it is finitely generated in degree $\leq C$ if there is a finite list of classes $\{v_{i}\in V_{d_{i}}\}$ with $d_{i}\leq C$, such that every $V_{n}$ is spanned by classes of the form $f^{*}v_{i}$.
Theorem 6.1 ([tosteson2021stability]).
Let $g,i\in\mathbb{N}$. Then the $\mathbf{FS}^{\text{op}}$-module
$n\mapsto H_{i}(\overline{\mathcal{M}}_{g,n};\mathbb{Q})$
is a subquotient of an extension of $\mathbf{FS}^{\text{op}}$-modules that are finitely generated in degree $\leq 8g^{2}i^{2}+29g^{2}i+16gi^{2}+21g^{2}+10gi-6g$.
Furthermore, Tosteson considers the functor $\text{Res}_{r}$ that restricts an $\mathbf{FS}^{\text{op}}$-module to an $\mathbf{FS}_{r}^{\text{op}}$-module, for $r\in\mathbb{N}$, where $\mathbf{FS}_{r}^{\text{op}}$ is the full subcategory of $\mathbf{FS}^{\text{op}}$ spanned by sets of size $\leq r$. The functor $\text{Res}_{r}$ has a left adjoint $\text{Ind}_{r}$, which takes an $\mathbf{FS}_{r}^{\text{op}}$ module to the $\mathbf{FS}^{\text{op}}$ module freely generated by it modulo relations in degree $r$.
Theorem 6.2 ([tosteson2021stability]).
Let $i,g\in\mathbb{N}$. There exists $N\in\mathbb{N}$ such that the natural map of $\mathbf{FS}^{\text{op}}$-modules
$\text{Ind}_{N}\text{Res}_{N}H_{i}(\overline{\mathcal{M}}_{g,-},\mathbb{Q})\to H_{i}(\overline{\mathcal{M}}_{g,-},\mathbb{Q})$
is an isomorphism. In particular, any presentation of the $\mathbf{FS}_{N}^{\text{op}}$-module $\text{Res}_{N}H_{i}(\overline{\mathcal{M}}_{g,-},\mathbb{Q})$ gives a presentation of the $\mathbf{FS}^{\text{op}}$-module $H_{i}(\overline{\mathcal{M}}_{g,-},\mathbb{Q})$.
@@ 6.1. Tosteson’s stability theorem in $n$
Our Theorem 5.6 recovers Tosteson’s stability theorems in $n$, noting that a Whitehead move (or WDVV exchange), on a stable graph preserves homology up to a nonzero factor. However, our bound on the stable range is weaker than Tosteson’s, in the case of fixed $g$, c.f. Remark 2.8.
6.2. Tautological extension
A tautological morphism is a morphism induced by gluing, forgetting, and permuting marked points. The tautological ring $RH^{*}(\overline{\mathcal{M}}_{g,n})\subset H^{*}(\overline{\mathcal{M}}_{g,n})$ is the smallest subring that is closed under all tautological morphisms.
Canning, Larson, and Payne [canning2024extensions] define a semi-tautological extension (STE) as a collection of subrings $S^{*}(\overline{\mathcal{M}}_{g,n})$ of $H^{*}(\overline{\mathcal{M}}_{g,n})$ that contains the tautological subrings $RH^{*}(\overline{\mathcal{M}}_{g,n})$ and is closed under pullback by forgetting and permuting marked points and under pushforward for gluing marked points. They prove
Theorem 6.4 ([canning2024extensions]).
For any fixed degree $k$, the STE generated by
$\{H^{k^{\prime}}(\overline{\mathcal{M}}_{g^{\prime},n^{\prime}}):k^{\prime}\leq k,g^{\prime}<\frac{3}{2}k^{\prime}+1,n^{\prime}\leq k^{\prime},4g^{\prime}-4+n^{\prime}\geq k\}$
contains $H^{k}(\overline{\mathcal{M}}_{g,n})$ for all $g$ and $n$.
In our setting, we only consider tautological morphisms of certain types, namely gluing (G) and permuting (P) marked points; forgetful maps are not included. Analogous to the perspective of Canning, Larson, and Payne, we may consider tautological extensions of type ‘GP’ in homology. Our Theorem 5.6 then says that a finitely generated tautological extension of type ‘GP’ contains $H_{k}(\overline{\mathcal{M}}_{g,n};\mathbb{Q})$ for all $(g,n)$.